### v1

# Stochastic theta methods for random periodic solution of stochastic differential equations under non-globally Lipschitz conditions11footnotemark: 1

Abstract
This work focuses on the numerical approximations of random periodic solutions of stochastic differential equations (SDEs). Under non-globally Lipschitz conditions, we prove the existence and uniqueness of random periodic solutions for the considered equations and its numerical approximations generated by the stochastic theta (ST) methods with $\theta\in(1/2,1]$. It is shown that the random periodic solution of each ST method converges strongly in the mean square sense to that of SDEs for all stepsize. More precisely, the mean square convergence order is $1/2$ for SDEs with multiplicative noise and $1$ for SDEs with additive noise. Numerical results are finally reported to confirm these theoretical findings.
AMS subject classification: 37H99, 60H10, 60H35, 65C30
Key Words: Stochastic theta methods, Random periodic solution, Multiplicative SDEs, mean square convergence order, non-globally Lipschitz condition
\$\$\$\$footnotetext: 11footnotemark: 1 This work was supported by National Natural Science Foundation of China (Nos. 12201552, 11961029 and 12031020), the Natural Science Foundation of Jiangxi Province (No. 2020BABL201007), Yunnan Fundamental Research Projects (No. 202301AU070010) and Innovation Team of School of Mathematics and Statistics of Yunnan University (No. ST20210104). \$\$\$\$footnotetext: 22footnotemark: 2 School of Mathematics and Statistics, Yunnan University, Kunming, Yunnan 650500, China. Email: czh@ynu.edu.cn\$\$\$\$footnotetext: 33footnotemark: 3 School of Mathematics and Statistics, Yunnan University, Kunming, Yunnan 650500, China. Email: caoliangmin@itc.ynu.edu.cn\$\$\$\$footnotetext: 44footnotemark: 4 School of Statistics and Data Science, Jiangxi University of Finance and Economics, Nanchang, Jiangxi 330013, China. Email: chenlin@jxufe.edu.cn. Corresponding author.
1 Introduction
As a very active ongoing research area, the study of numerical solutions of SDEs has achieved a large amount of mathematical results in the past decades; see, e.g., monographs [13, 17] for SDEs with the traditional global Lipschitz condition and [8, 11, 12, 20, 19, 9, 15, 16, 2, 1, 10, 4, 21, 22, 7] for SDEs beyond such a restrictive condition. However, just a very limited number of literature [4, 21, 22, 7] began to study the numerical approximation of random periodic solutions of SDEs since the concept of random periodic solutions of random dynamical systems has been properly defined in [23]. The periodicity of random periodic solutions of SDEs means that there exists a random invariant curve with period $\tau>0$, which will be mapped into itself with respect to the shifted noise. In general, the random invariant curves of random periodic solutions are rarely available in explicit form. Therefore it is natural to take advantage of numerical methods to characterize the random phenomena with a periodic nature.
We consider the following semi-linear SDE with multiplicative noise
$\text{d}X_{t}^{t_{0}}=\big{(}-AX_{t}^{t_{0}}+f(t,X_{t}^{t_{0}})\big{)}{\,\rm{d}}{t}+g(t,X^{t_{0}}_{t}){\,\rm{d}}{W_{t}},\quad t\geq t_{0}$ (1.1)
with $X_{t_{0}}^{t_{0}}=\xi$, where $f\colon\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, $g\colon\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are continuous functions and $A$ is a positive definite $d\times d$ symmetric matrix. Besides, $\{W_{t}\}_{t\in\mathbb{R}}$ is a standard two-sided Wiener process in $\mathbb{R}^{m}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. Here the notation $X_{t}^{t_{0}},t\geq t_{0}$ denotes the solution process starting from $t_{0}$. Under Assumption 2.1 specified later, (1.1) admits a unique solution $\{X_{t}^{t_{0}}\}_{t\geq t_{0}}$ (we also rewrite it as $\{X_{t}^{t_{0}}(\xi)\}_{t\geq t_{0}}$ to emphasize its dependence on the initial value $\xi$), described by the variation of constant formula as follows
$X_{t}^{t_{0}}(\xi)=e^{-A(t-t_{0})}\xi+\int_{t_{0}}^{t}e^{-A(t-s)}f(s,X_{s}^{t_{0}}){\,\rm{d}}{s}+\int_{t_{0}}^{t}e^{-A(t-s)}g(s,X_{s}^{t_{0}}){\,\rm{d}}{W_{s}},\quad t\geq t_{0},$ (1.2)
Then the map $u(t,t_{0})\colon\Omega\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined by $u(t,t_{0})\xi=X_{t}^{t_{0}}(\xi)$ satisfies the semi-flow property $u(t,r,\omega)=u(t,s,\omega)\circ u(s,r,\omega)$ and the periodic property (with period $\tau>0$) $u(t+\tau,s+\tau,\omega)=u(t,s,\Theta_{t}\omega)$ with $\Theta_{t}(\omega)(s):=W_{t+s}-W_{t}$ for all $s\leq r\leq t,\omega\in\Omega$. Theorem 2.4 shows that the pull-back sequence $\{X_{t}^{-k\tau}(\xi)\}_{k\in\mathbb{N}}$ has a limit $X_{t}^{*}$ in $L^{2}(\Omega;\mathbb{R}^{d})$, where $X_{t}^{*}$ is the random periodic solution of (1.1) and is given by
$X_{t}^{*}=\int_{-\infty}^{t}e^{-A(t-s)}f(s,X_{s}^{*}){\,\rm{d}}{s}+\int_{-\infty}^{t}e^{-A(t-s)}g(s,X_{s}^{*}){\,\rm{d}}{W_{s}}.$ (1.3)
Then the numerical approximation of the random periodic solution $X_{t}^{*}$ becomes a numerical approximation of an infinite time horizon problem.
Concerning the numerical approximation of random periodic solution, the seminal work [4] establishes the standard mean square convergence orders of the Euler–Maruyama method and a modified Milstein method for dissipative SDEs with multiplicative noise under the global Lipschitz conditions. Subsequently, [21] generalizes the Euler–Maruyama method to the ST methods and derives the convergence order $1/4$. Taking into account that most SDEs arising from applications possess super-linearly or sub-linearly growing coefficients, it is necessary to explore the numerical results for SDEs beyond the global Lipschitz conditions. For SDEs with additive noise under the one-sided Lipschitz condition, [22] proves that the random periodic solution of the backward Euler–Maruyama method converges strongly in the mean square sense to that of SDEs with order $1/2$, which has been lifted to order $1$ under a more relaxed condition in [7]. However, to the best of our knowledge, there are not any results about the numerical approximation of random periodic solution for SDEs with multiplicative noise under non-globally Lipschitz conditions, which motivates us to make a contribution to this problem.
Noting that we are dealing with a numerical approximation of an infinite time horizon problem, we take a stepsize $\Delta\in(0,1)$ and define an equidistant partition $\mathcal{P}^{\Delta}:=\{t_{j}:=j\Delta\}_{j\in\mathbb{Z}}$ for the real line $\mathbb{R}$. Throughout this work, we always let the parameter $\theta\in(1/2,1]$. In order to numerically approximate the solution process $\{X_{t}^{-k\tau}\}_{t\geq-k\tau}$ of (1.1) starting at $-k\tau$ with $k\in\mathbb{N}$, the ST methods are given by
$\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}=$ $~{}\hat{X}_{-k\tau+t_{j}}^{-k\tau}+\theta\Delta\big{(}-A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})\big{)}$
$~{}+(1-\theta)\Delta\big{(}-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})\big{)}+g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})\Delta W_{-k\tau+t_{j}}$ (1.4)
for all $j\in N$, where $\hat{X}_{-k\tau}^{-k\tau}=\xi$ is the initial value and $\Delta W_{-k\tau+t_{j}}:=W_{-k\tau+t_{j+1}}-W_{-k\tau+t_{j}}$ is the Brownian increment. Based on the uniform bounds for the second moment of the numerical approximation $\{\hat{X}_{-k\tau+t_{j}}^{-k\tau}\}_{j\in N}$, we employ Lemma 3.2 to show that each ST method admits a unique random periodic solution $\hat{X}_{t}^{*}$. Without resorting to the discrete Gronwall inequality, a technical iterative argument is introduced in the error estimate to overcome the difficulties arising by the superlinear growth of coefficients and the unboundedness of the time intervals. With the help of the Hölder continuity of the exact solution, we derive a uniform bound for the error between $\hat{X}_{-k\tau+t_{j}}^{-k\tau}$ and $X_{-k\tau+t_{j}}^{-k\tau}$ for all $k,j\in\mathbb{N}$, and consequently establish the mean square convergence of random periodic solutions. To be precise, the mean square convergence order of each ST method is $1/2$ for SDEs with multiplicative noise and $1$ for SDEs with additive noise, which, in particular, coincides with [7, Theorem 4.6] investigating ST method with $\theta=1$ for SDEs with additive noise.
The remainder of this paper is organized as follows. Section 2 contains some assumptions and the existence and uniqueness of random periodic solutions of SDEs (1.1). In Section 3, we present the well-posedness and the existence of unique random periodic solutions of the ST methods. The mean square convergence for random periodic solutions of the ST methods is established in Section 4. Finally, several numerical experiments are given in Section 5 to illustrate the theoretical results.
2 Assumptions and random periodic solutions of SDEs
[…]
In order to provide efficient numerical approximation for SDEs (1.1), we put the following conditions on (1.1).
Assumption 2.1.
Suppose the following conditions are satisfied. (A1) The eigenvalues $\{\lambda_{i}\}_{i=1,2,\cdots,d}$ of the symmetric matrix $A$ satisfy $\lambda:=\min\{\lambda_{i}\}_{i=1,2,\cdots,d}>0.$ (A2) The coefficients $f$ and $g$ are continuous and periodic in time with period $\tau>0$, i.e.,
$f(t+\tau,x)=f(t,x),\quad g(t+\tau,x)=g(t,x),\quad x\in\mathbb{R}^{d},~{}t\in\mathbb{R}.$ (2.1)
(A3) There exist constants $\gamma\geq 1$, $p^{*}>4\gamma-2$, $L_{f}\in(0,\lambda)$ and $C>0$ such that for all $x,y\in\mathbb{R}^{d}$ and $t,s\in\mathbb{R}$,
$\begin{split}&~{}\langle x-y,f(t,x)-f(s,y)\rangle+(p^{*}-1)|g(t,x)-g(s,y)|^{2}\\ \leq&~{}L_{f}|x-y|^{2}+C(1+|x|+|y|)^{\gamma+1}|t-s|,\quad x,y\in\mathbb{R}^{d},t,s\in\mathbb{R},\end{split}$ (2.2)
and
$|f(t,x)-f(s,y)|\leq C\big{(}(1+|x|+|y|)^{\gamma-1}|x-y|+(1+|x|+|y|)^{\gamma}|t-s|\big{)}.$ (2.3)
(A4) For any $p>0$, there exists a constant $C>0$ such that $|\xi|_{L^{p}(\Omega;\mathbb{R}^{d})}\leq C$.
For later convenience, we note several consequences of Assumption 2.1. (A1) implies
$\big{\langle}-Ax,x\big{\rangle}\leq-\lambda|x|^{2},\quad x\in\mathbb{R}^{d}.$ (2.4)
From (2.2), we obtain
$\langle x-y,f(t,x)-f(t,y)\rangle+(p^{*}-1)|g(t,x)-g(t,y)|^{2}\leq L_{f}|x-y|^{2},\quad x,y\in\mathbb{R}^{d},t\in\mathbb{R},$ (2.5)
which together with the weighted Young inequality shows that for any $x\in\mathbb{R}^{d},t\in\mathbb{R}$,
$~{}\langle x,f(t,x)\rangle+\frac{p^{*}-1}{2}|g(t,x)|^{2}$
$\leq$ $~{}\langle x,f(t,x)-f(t,0)\rangle+(p^{*}-1)|g(t,x)-g(t,0)|^{2}+\langle x,f(t,0)\rangle+(p^{*}-1)|g(t,0)|^{2}$
$\leq$ $~{}L_{f}|x|^{2}+\frac{\lambda-L_{f}}{2}|x|^{2}+\frac{1}{2(\lambda-L_{f})}|f(t,0)|^{2}+(p^{*}-1)|g(t,0)|^{2}$ (2.6)
$\leq$ $~{}\widetilde{L_{f}}|x|^{2}+\widetilde{C}$
with
$\widetilde{L_{f}}:=\frac{\lambda+L_{f}}{2}\in\big{(}L_{f},\lambda\big{)},\quad\widetilde{C}:=\frac{1}{2(\lambda-L_{f})}\bigg{(}\sup_{t\in[0,\tau)}|f(t,0)|^{2}\bigg{)}+(p^{*}-1)\bigg{(}\sup_{t\in[0,\tau)}|g(t,0)|^{2}\bigg{)}>0.$
Besides, combining (2.2) and (2.3) shows that there exists a constant $C>0$ such that
$|g(t,x)-g(s,y)|^{2}\leq C(1+|x|+|y|)^{\gamma-1}|x-y|^{2}+C(1+|x|+|y|)^{\gamma+1}|t-s|$ (2.7)
for all $x,y\in\mathbb{R}^{d}$ and $t,s\in\mathbb{R}$. Together with (2.3), we derive
$|f(t,x)|\leq C(1+|x|)^{\gamma},\quad|g(t,x)|^{2}\leq C(1+|x|)^{\gamma+1},\quad x\in\mathbb{R}^{d},t\in\mathbb{R}.$ (2.8)
According to (2.3) and (2.7), we know that $f$ and $g$ are locally Lipschitz continuous. In combination with the monotone condition (2), [14, Theorem 3.5 in Chapter 2] ensures that (1.1) admits a unique strong solution $\{X_{t}^{t_{0}}\}_{t\geq t_{0}}$, given by
$X_{t}^{t_{0}}=\xi+\int_{t_{0}}^{t}-AX_{s}^{t_{0}}+f(s,X_{s}^{t_{0}}){\,\rm{d}}{s}+\int_{t_{0}}^{t}g(s,X_{s}^{t_{0}}){\,\rm{d}}{W_{s}},\quad t\geq t_{0}.$ (2.9)
To establish the existence and uniqueness of random periodic solution of (1.1), two auxiliary lemmas are stated and proved below. The first one considers the uniform boundedness for the exact solution of (1.1).
Lemma 2.2.
Suppose that Assumption 2.1 holds. Then for any $p\in[2,p^{*}]$, there exists a constant $C>0$ such that
$\sup_{k\in\mathbb{N}}\sup_{t\geq-k\tau}\mathbb{E}\big{[}|X_{t}^{-k\tau}|^{p}\big{]}\leq C.$ (2.10)
Proof.
Applying Itô formula shows that for any $k\in\mathbb{N}$, $t\geq-k\tau$ and $p\in(2,p^{*}]$,
$e^{p\lambda t}|X_{t}^{-k\tau}|^{p}=$ $~{}e^{-p\lambda k\tau}|\xi|^{p}+\int_{-k\tau}^{t}p\lambda e^{p\lambda s}|X_{s}^{-k\tau}|^{p}+pe^{p\lambda s}|X_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau},-AX_{s}^{-k\tau}\big{\rangle}$
$~{}+pe^{p\lambda s}|X_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau},f(s,X_{s}^{-k\tau})\big{\rangle}+\frac{p}{2}e^{p\lambda s}|X_{s}^{-k\tau}|^{p-2}|g(s,X_{s}^{-k\tau})|^{2}$
$~{}+\frac{p(p-2)}{2}e^{p\lambda s}|X_{s}^{-k\tau}|^{p-4}|(X_{s}^{-k\tau})^{\top}g(s,X_{s}^{-k\tau})|^{2}{\,\rm{d}}{s}$
$~{}+\int_{-k\tau}^{t}pe^{p\lambda s}|X_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau},g(s,X_{s}^{-k\tau}){\,\rm{d}}{W_{s}}\big{\rangle}.$
It follows from the Cauchy–Schwarz inequality, (2.4) and (2) that
$~{}\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}|^{p}\big{]}$
$\leq$ $~{}\mathbb{E}\big{[}e^{-p\lambda k\tau}|\xi|^{p}\big{]}+\int_{-k\tau}^{t}pe^{p\lambda s}\mathbb{E}\big{[}|X_{s}^{-k\tau}|^{p-2}\big{(}\lambda|X_{s}^{-k\tau}|^{2}+\big{\langle}X_{s}^{-k\tau},-AX_{s}^{-k\tau}\big{\rangle}\big{)}\big{]}{\,\rm{d}}{s}$
$~{}+\int_{-k\tau}^{t}pe^{p\lambda s}\mathbb{E}\Big{[}|X_{s}^{-k\tau}|^{p-2}\Big{(}\big{\langle}X_{s}^{-k\tau},f(s,X_{s}^{-k\tau})\big{\rangle}+\frac{p-1}{2}|g(s,X_{s}^{-k\tau})|^{2}\Big{)}\Big{]}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}|\xi|^{p}\big{]}+\int_{-k\tau}^{t}pe^{p\lambda s}\big{(}\widetilde{L_{f}}\mathbb{E}\big{[}|X_{s}^{-k\tau}|^{p}\big{]}+\widetilde{C}\mathbb{E}\big{[}|X_{s}^{-k\tau}|^{p-2}\big{]}\big{)}{\,\rm{d}}{s}.$
Due to the Young inequality $|x|^{p-2}|y|\leq\frac{p-2}{p}|x|^{p}+\frac{2}{p}|y|^{\frac{p}{2}}$ for any $x,y\in\mathbb{R}$, we choose a fixed constant $\varepsilon\in(0,\lambda-\widetilde{L_{f}})$ to derive
$\widetilde{C}|X_{s}^{-k\tau}|^{p-2}=$ $~{}\frac{p}{p-2}\times\bigg{(}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{p-2}{p}}|X_{s}^{-k\tau}|^{p-2}\times\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\bigg{)}$
$\leq$ $~{}\frac{p}{p-2}\times\bigg{(}\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)|X_{s}^{-k\tau}|^{p}+\frac{2}{p}\Big{(}\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\Big{)}^{\frac{p}{2}}\bigg{)}$
$\leq$ $~{}(\lambda-\widetilde{L_{f}}-\varepsilon)|X_{s}^{-k\tau}|^{p}+\frac{2}{p-2}\big{(}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\big{)}^{\frac{p}{2}},$
and consequently
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}|^{p}\big{]}\leq\mathbb{E}\big{[}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\int_{-k\tau}^{t}p(\lambda-\varepsilon)\mathbb{E}\big{[}e^{p\lambda s}|X_{s}^{-k\tau}|^{p}\big{]}{\,\rm{d}}{s}.$
The Gronwall inequality [3, Theorem 1] implies
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\int_{-k\tau}^{t}p(\lambda-\varepsilon)\big{(}\mathbb{E}\big{[}|\xi|^{p}\big{]}+Ce^{p\lambda s}\big{)}e^{\int_{s}^{t}p(\lambda-\varepsilon){\,\rm{d}}{u}}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\mathbb{E}\big{[}|\xi|^{p}\big{]}e^{p(\lambda-\varepsilon)(t+k\tau)}+C\frac{\lambda-\varepsilon}{\varepsilon}e^{p\lambda t}.$
Then we obtain (2.10) for $p\in(2,p^{*}]$, which together with the Hölder inequality indicates (2.10) for $p=2$. Thus we finish the proof. ∎
The forthcoming lemma explores the dependence of solutions of (1.1) on different initial values.
Lemma 2.3.
Let $\{X_{t}^{-k\tau}\}_{t\geq-k\tau}$ and $\{Y_{t}^{-k\tau}\}_{t\geq-k\tau}$ be two solutions of (1.1) with different initial values $\xi$ and $\eta$, respectively. Suppose that Assumption 2.1 holds for both initial values $\xi$ and $\eta$. Then for any $p\in[2,p^{*}],k\in\mathbb{N}$ and $t\geq-k\tau$,
$\mathbb{E}\big{[}|X_{t}^{-k\tau}-Y_{t}^{-k\tau}|^{p}\big{]}\leq e^{p(L_{f}-\lambda)(t+k\tau)}\mathbb{E}\big{[}|\xi-\eta|^{p}\big{]}.$ (2.11)
Proof.
Because of $X_{-k\tau}^{-k\tau}-Y_{-k\tau}^{-k\tau}=\xi-\eta$ and
${\,\rm{d}}{(X_{t}^{-k\tau}-Y_{t}^{-k\tau})}=$ $~{}\big{(}-A(X_{t}^{-k\tau}-Y_{t}^{-k\tau})+(f(t,X_{t}^{-k\tau})-f(t,Y_{t}^{-k\tau}))\big{)}{\,\rm{d}}{t}$
$~{}+\big{(}g(t,X_{t}^{-k\tau})-g(t,Y_{t}^{-k\tau})\big{)}{\,\rm{d}}{W_{t}},\quad t>-k\tau,$
we use the Itô formula to get
$~{}e^{p\lambda t}|X_{t}^{-k\tau}-Y_{t}^{-k\tau}|^{p}$
$=$ $~{}e^{-p\lambda k\tau}|\xi-\eta|^{p}+\int_{-k\tau}^{t}p\lambda e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p}$
$~{}+pe^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau}-Y_{s}^{-k\tau},-A(X_{s}^{-k\tau}-Y_{s}^{-k\tau})\big{\rangle}{\,\rm{d}}{s}$
$~{}+\int_{-k\tau}^{t}pe^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau}-Y_{s}^{-k\tau},f(s,X_{s}^{-k\tau})-f(s,Y_{s}^{-k\tau})\big{\rangle}$
$~{}+\frac{p}{2}e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}|g(s,X_{s}^{-k\tau})-g(s,Y_{s}^{-k\tau})|^{2}$
$~{}+\frac{p(p-2)}{2}e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-4}|(X_{s}^{-k\tau}-Y_{s}^{-k\tau})^{\top}g(s,X_{s}^{-k\tau})-g(s,Y_{s}^{-k\tau})|^{2}{\,\rm{d}}{s}$
$~{}+\int_{-k\tau}^{t}pe^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau}-Y_{s}^{-k\tau},(g(s,X_{s}^{-k\tau})-g(s,Y_{s}^{-k\tau})){\,\rm{d}}{W_{s}}\big{\rangle}{\,\rm{d}}{s}.$
Taking expectations and utilizing the Cauchy–Schwarz inequality, (2.2) as well as (2.4) show that for any $p\in[2,p^{*}]$,
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}-Y_{t}^{-k\tau}|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}e^{-p\lambda k\tau}|\xi-\eta|^{p}\big{]}+\int_{-k\tau}^{t}p\mathbb{E}\Big{[}e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}$
×$~{}\times\Big{(}\big{\langle}X_{s}^{-k\tau}-Y_{s}^{-k\tau},f(s,X_{s}^{-k\tau})-f(s,Y_{s}^{-k\tau})\big{\rangle}$
$~{}+\frac{p-1}{2}|g(s,X_{s}^{-k\tau})-g(s,Y_{s}^{-k\tau})|^{2}\Big{)}\Big{]}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}e^{-p\lambda k\tau}|\xi-\eta|^{p}\big{]}+\int_{-k\tau}^{t}pL_{f}\mathbb{E}\big{[}e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p}\big{]}{\,\rm{d}}{s}.$
It follows from the Gronwall inequality [3, Corollary 3]that
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}-Y_{t}^{-k\tau}|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}e^{-p\lambda k\tau}|\xi-\eta|^{p}\big{]}e^{pL_{f}(t+k\tau)},$
which immediately results in (2.11). ∎
Now we can prove the existence and uniqueness of random periodic solution for (1.1), whose proof is analogous to that of [4, Theorem 2.4] and hence is omitted.
Theorem 2.4.
Suppose that Assumption 2.1 holds. Then for any initial value $\xi$ satisfying (A4), the considered equation (1.1) admits a unique random periodic solution $X_{t}^{*}(\cdot)\in L^{2}(\Omega),t\geq 0$, given by
$X_{t}^{*}=\int_{-\infty}^{t}e^{-A(t-s)}f(s,X_{s}^{*}){\,\rm{d}}{s}+\int_{-\infty}^{t}e^{-A(t-s)}g(s,X_{s}^{*}){\,\rm{d}}{W_{s}},\quad t\geq 0$
such that
$\lim_{k\rightarrow\infty}|X_{t}^{-k\tau}-X_{t}^{*}|_{L^{2}(\Omega;\mathbb{R}^{d})}=0.$ (2.12)
3 Random periodic solutions of ST methods
This section aims to show the existence and uniqueness of random periodic solutions for the ST methods. For this purpose, let us first verify the well-posedness of (1). In fact, (2.5) implies
$\langle x-y,f(t,x)-f(t,y)\rangle\leq L_{f}|x-y|^{2},\quad x,y\in\mathbb{R}^{d},t\in\mathbb{R},$
According to the uniform monotonicity theorem [18, Theorem C.2] and using $L_{f}\in(0,\lambda)$, we know that (1) admits a unique solution $\{\hat{X}_{-k\tau+t_{j}}^{-k\tau}\}_{j\in\mathbb{N}}$ for any stepsize $\Delta\in(0,1)$. Besides, there is a uniform bound for the second moment of the numerical solution $\{\hat{X}_{-k\tau+t_{j}}^{-k\tau}\}_{j\in\mathbb{N}}$, as indicated below.
Lemma 3.1.
Suppose that Assumption 2.1 holds and that $\theta\in(\frac{1}{2},1]$. Then there exists a constant $C>0$, independent of $j,k$ and $\Delta$, such that
$\sup_{k,j\in\mathbb{N}}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}\leq C.$ (3.1)
Proof.
Noticing that (1) implies
$~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}-\theta\Delta\big{(}-A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})\big{)}|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})\big{)}+g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})\Delta W_{-k\tau+t_{j}}|^{2}\big{]}$ (3.2)
$=$ $~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})\big{)}|^{2}\big{]}+\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]},$
where we have used
$\mathbb{E}\big{[}\big{\langle}\hat{X}_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau}),g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})\Delta{W_{-k\tau+t_{j}}}\big{\rangle}\big{]}=0$
due to the $\mathcal{F}_{-k\tau+t_{j}}$-measurability of $\hat{X}_{-k\tau+t_{j}}^{-k\tau}$ and the independence between $\hat{X}_{-k\tau+t_{j}}^{-k\tau}$ and $\Delta{W_{-k\tau+t_{j}}}$ for all $j$. On one hand, (2.4) and (2) enable us to show
$~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}-\theta\Delta\big{(}-A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})\big{)}|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}-f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+2\theta\Delta\mathbb{E}\big{[}\big{\langle}\hat{X}_{-k\tau+t_{j+1}}^{-k\tau},A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}\big{\rangle}\big{]}-2\theta\Delta\mathbb{E}\big{[}\big{\langle}\hat{X}_{-k\tau+t_{j+1}}^{-k\tau},f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})\big{\rangle}\big{]}$
$\geq$ $~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}-f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+2\lambda\theta\Delta\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}-2\widetilde{L_{f}}\theta\Delta\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}-2\widetilde{C}\theta\Delta$
$=$ $~{}\big{(}1+2(\lambda-\widetilde{L_{f}})\theta\Delta\big{)}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}-f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}-2\widetilde{C}\theta\Delta.$
On the other hand, using (2.4) and (2) again yields
$~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})\big{)}|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+2(1-\theta)\Delta\mathbb{E}\big{[}\big{\langle}\hat{X}_{-k\tau+t_{j}}^{-k\tau},-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})\big{\rangle}\big{]}$
$\leq$ $~{}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}-2\lambda(1-\theta)\Delta\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+2\widetilde{L_{f}}(1-\theta)\Delta\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+2(1-\theta)\Delta\widetilde{C}-(p^{*}-1)(1-\theta)\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$=$ $~{}\big{(}1-2(\lambda-\widetilde{L_{f}})(1-\theta)\Delta\big{)}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+2(1-\theta)\Delta\widetilde{C}-(p^{*}-1)(1-\theta)\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}.$
It follows that
$~{}\big{(}1+2(\lambda-\widetilde{L_{f}})\theta\Delta\big{)}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}-f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$\leq$ $~{}\big{(}1-2(\lambda-\widetilde{L_{f}})(1-\theta)\Delta\big{)}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+\big{(}(p^{*}-1)\theta-(p^{*}-2)\big{)}\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}+2\Delta\widetilde{C}$
$\leq$ $~{}\big{(}1+2(\lambda-\widetilde{L_{f}})\theta\Delta-2(\lambda-\widetilde{L_{f}})\theta\Delta\big{)}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+\Big{(}(p^{*}-1)\theta-\frac{p^{*}-2}{2}\Big{)}\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+(\theta^{2}-2\theta+1)\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}+2\Delta\widetilde{C}.$
Setting
$\widetilde{C_{\Delta}}:=\max\bigg{\{}1-\frac{2(\lambda-\widetilde{L_{f}})\theta\Delta}{1+2(\lambda-\widetilde{L_{f}})\theta\Delta},~{}1-\frac{p^{*}-2}{2(p^{*}-1)\theta},~{}1-\frac{2\theta-1}{\theta^{2}}\bigg{\}}\in[0,1)$
leads to
$~{}\big{(}1+2(\lambda-\widetilde{L_{f}})\theta\Delta\big{)}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|A\hat{X}_{-k\tau+t_{j+1}}^{-k\tau}-f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}-\frac{2\Delta\widetilde{C}}{1-\widetilde{C_{\Delta}}}$
$\leq$ $~{}\widetilde{C_{\Delta}}\Big{(}\big{(}1+2(\lambda-\widetilde{L_{f}})\theta\Delta\big{)}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-A\hat{X}_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}-\frac{2\Delta\widetilde{C}}{1-\widetilde{C_{\Delta}}}\Big{)}.$
By iteration, we use (2.8) to obtain
$~{}\big{(}1+2(\lambda-\widetilde{L_{f}})\theta\Delta\big{)}\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}-\frac{2\Delta\widetilde{C}}{1-\widetilde{C_{\Delta}}}$
$\leq$ $~{}\widetilde{C_{\Delta}}^{j}\bigg{(}\big{(}1+2(\lambda-\widetilde{L_{f}})\theta\Delta\big{)}\mathbb{E}\big{[}|\xi|^{2}\big{]}+(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(0,\xi)|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-A\xi+f(0,\xi)|^{2}\big{]}-\frac{2\Delta\widetilde{C}}{1-\widetilde{C_{\Delta}}}\bigg{)}$
$\leq$ $~{}C\big{(}1+\mathbb{E}\big{[}|\xi|^{2}\big{]}+\mathbb{E}\big{[}|\xi|^{\gamma+1}\big{]}+\mathbb{E}\big{[}|\xi|^{2\gamma}\big{]}\big{)}.$
Together with
$\frac{\Delta}{1-\widetilde{C_{\Delta}}}=$ $~{}\Delta\max\bigg{\{}\frac{1+2(\lambda-\widetilde{L_{f}})\theta\Delta}{2(\lambda-\widetilde{L_{f}})\theta\Delta},~{}\frac{2(p^{*}-1)\theta}{p^{*}-2},~{}\frac{\theta^{2}}{2\theta-1}\bigg{\}}$
$=$ $~{}\max\bigg{\{}\frac{1+2(\lambda-\widetilde{L_{f}})\theta\Delta}{2(\lambda-\widetilde{L_{f}})\theta},~{}\frac{2(p^{*}-1)\theta\Delta}{p^{*}-2},~{}\frac{\theta^{2}\Delta}{2\theta-1}\bigg{\}}$
$\leq$ $~{}\max\bigg{\{}\frac{1+2(\lambda-\widetilde{L_{f}})\theta}{2(\lambda-\widetilde{L_{f}})\theta},~{}\frac{(p^{*}-1)\theta}{p^{*}-2},~{}\frac{\theta^{2}}{2\theta-1}\bigg{\}}:=\bar{C},$
we derive
$\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}\leq C\big{(}1+\mathbb{E}\big{[}|\xi|^{2}\big{]}+\mathbb{E}\big{[}|\xi|^{\gamma+1}\big{]}+\mathbb{E}\big{[}|\xi|^{2\gamma}\big{]}\big{)}+2\bar{C}\widetilde{C},$
and finish the proof. ∎
The following lemma indicates that any two numerical solutions starting from different initial conditions can be arbitrarily close after sufficiently many iterations.
Lemma 3.2.
Let $\{\hat{X}_{-k\tau+t_{j}}^{-k\tau}\}$ and $\{\hat{Y}_{-k\tau+t_{j}}^{-k\tau}\}$ be two solutions of the BEM scheme (1) with different initial values $\xi$ and $\eta$, respectively. Suppose that Assumption 2.1 holds for both initial values $\xi$ and $\eta$. Then there exist constants $C>0$ and $C_{\Delta}\in[0,1)$, independent of $j$ and $k$, such that
$\mathbb{E}\big{[}|\hat{X}_{-k\tau+t_{j}}^{-k\tau}-\hat{Y}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}\leq CC_{\Delta}^{j}\big{(}\mathbb{E}[|\xi-\eta|^{2}]+\mathbb{E}[|\xi-\eta|^{\gamma}]\big{)},$ (3.3)
where
$C_{\Delta}:=\max\bigg{\{}1-\frac{2(\lambda-L_{f})\theta\Delta}{1+2(\lambda-L_{f})\theta\Delta},~{}1-\frac{p^{*}-2}{2(p^{*}-1)\theta},~{}1-\frac{2\theta-1}{\theta^{2}}\bigg{\}}\in[0,1).$ (3.4)
Proof.
To simplify the notation, we denote $E_{-k\tau+t_{j}}^{-k\tau}:=\hat{X}_{-k\tau+t_{j}}^{-k\tau}-\hat{Y}_{-k\tau+t_{j}}^{-k\tau}$ and apply (1) as well as the arguments used in (3) to get
$~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j+1}}^{-k\tau}-\theta\Delta(-AE_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-f(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau}))|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta(-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau}))$
$~{}+(g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-g(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau}))\Delta W_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$ (3.5)
$=$ $~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta(-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau}))|^{2}\big{]}$
$~{}+\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-g(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}.$
It follows from (2.2) and (2.4) that
$~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j+1}}^{-k\tau}-\theta\Delta(-AE_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-f(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau}))|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-f(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}-2\theta\Delta\mathbb{E}\big{[}\big{\langle}E_{-k\tau+t_{j+1}}^{-k\tau},-AE_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-f(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau})\big{\rangle}\big{]}$ (3.6)
$\geq$ $~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|E_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-f(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-g(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}.$
Similarly, we have
$~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta(-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau}))|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|(-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau}))|^{2}\big{]}$
$~{}+2(1-\theta)\Delta\mathbb{E}\big{[}\big{\langle}E_{-k\tau+t_{j}}^{-k\tau},-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})\big{\rangle}\big{]}$
$\leq$ $~{}\big{(}1-2(\lambda-L_{f})(1-\theta)\Delta\big{)}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$ (3.7)
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}-2(p^{*}-1)(1-\theta)\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-g(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}.$
Inserting (3) and (3) into (3) results in
$~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|E_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-f(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-g(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$\leq$ $~{}\big{(}1-2(\lambda-L_{f})(1-\theta)\Delta\big{)}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+\big{(}1-2(p^{*}-1)(1-\theta)\big{)}\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-g(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$\leq$ $~{}\big{(}1+2(\lambda-L_{f})\theta\Delta-2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+(\theta^{2}-2\theta+1)\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+\big{(}2(p^{*}-1)\theta-(p^{*}-2)\big{)}\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-g(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]},$
which together with (3.4) implies
$~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|E_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j+1}}^{-k\tau}+f(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-f(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j+1},\hat{X}_{-k\tau+t_{j+1}}^{-k\tau})-g(t_{j+1},\hat{Y}_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$\leq$ $~{}C_{\Delta}\Big{(}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-AE_{-k\tau+t_{j}}^{-k\tau}+f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-f(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})-g(t_{j},\hat{Y}_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}\Big{)}.$
By (2.3), (2.7) and the Hölder inequality, one gets
$~{}\mathbb{E}\big{[}|E_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$\leq$ $~{}C_{\Delta}^{j}\Big{(}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|\xi-\eta|^{2}\big{]}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g(0,\xi)-g(0,\eta)|^{2}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|-A(\xi-\eta)+f(0,\xi)-f(0,\eta)|^{2}\big{]}\Big{)}$
$\leq$ $~{}CC_{\Delta}^{j}\big{(}\mathbb{E}\big{[}|\xi-\eta|^{2}\big{]}+\mathbb{E}\big{[}|f(0,\xi)-f(0,\eta)|^{2}\big{]}+\mathbb{E}\big{[}|g(0,\xi)-g(0,\eta)|^{2}\big{]}\big{)}$
$\leq$ $~{}CC_{\Delta}^{j}\big{(}\mathbb{E}\big{[}|\xi-\eta|^{2}\big{]}+\mathbb{E}\big{[}(1+|\xi|+|\eta|)^{2\gamma-2}|\xi-\eta|^{2}\big{]}+\mathbb{E}\big{[}(1+|\xi|+|\eta|)^{\gamma-1}|\xi-\eta|^{2}\big{]}\big{)}$
$\leq$ $~{}CC_{\Delta}^{j}\Big{(}\mathbb{E}\big{[}|\xi-\eta|^{2}\big{]}+\big{(}\mathbb{E}\big{[}(1+|\xi|+|\eta|)^{2\gamma}\big{]}\big{)}^{\frac{\gamma-1}{\gamma}}\big{(}\mathbb{E}\big{[}|\xi-\eta|^{2\gamma}\big{]}\big{)}^{\frac{1}{\gamma}}$
$~{}+\big{(}\mathbb{E}\big{[}(1+|\xi|+|\eta|)^{\gamma}\big{]}\big{)}^{\frac{\gamma-1}{\gamma}}\big{(}\mathbb{E}\big{[}|\xi-\eta|^{2\gamma}\big{]}\big{)}^{\frac{1}{\gamma}}\Big{)},$
$\leq$ $~{}CC_{\Delta}^{j}\Big{(}\mathbb{E}\big{[}|\xi-\eta|^{2}\big{]}+\big{(}\mathbb{E}\big{[}|\xi-\eta|^{2\gamma}\big{]}\big{)}^{\frac{1}{\gamma}}\Big{)},$
which indicates the required result and therefore completes the proof. ∎
Following the arguments used in [4, Theorem 3.4], [21, Theorem 2.1] and [22, Theorem 8], we obtain the existence and uniqueness of random periodic solutions for the ST methods.
Theorem 3.3.
Suppose that Assumption 2.1 holds. Then for any initial value $\xi$ satisfying (A4), the considered method (1) admits a unique random period solution $\hat{X}_{t}^{*}(\cdot)\in L^{2}(\Omega;\mathbb{R}^{d}),t\geq 0$ such that
$\lim_{k\to\infty}|\hat{X}_{t}^{-k\tau}-\hat{X}_{t}^{*}|_{L^{2}(\Omega;\mathbb{R}^{d})}=0.$ (3.8)
4 Mean square convergence order of ST methods
In this section, we will establish the mean square convergence order of ST methods. To this end, the Hölder continuity of the exact solution is needed.
Lemma 4.1.
Suppose that Assumption 2.1 holds. Then for any $q\in[2,\frac{p^{*}}{\gamma}]$, $k\in\mathbb{N}$ and $s,t\geq-k\tau$, there exists a constant $C>0$, independent of $k,s,t$, such that
$\mathbb{E}\big{[}|X_{t}^{-k\tau}-X_{s}^{-k\tau}|^{q}\big{]}\leq C\big{(}|t-s|^{\frac{q}{2}}+|t-s|^{q}\big{)}.$ (4.1)
Proof.
Without loss of generality, we assume $s\leq t$ and get
$~{}\mathbb{E}\big{[}|X_{t}^{-k\tau}-X_{s}^{-k\tau}|^{q}\big{]}$
$=$ $~{}\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{-k\tau}+f(r,X_{r}^{-k\tau}){\,\rm{d}}{r}+\int_{s}^{t}g(r,X_{r}^{-k\tau}){\,\rm{d}}{W_{r}}\Big{|}^{q}\bigg{]}$ (4.2)
$\leq$ $~{}C\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{-k\tau}+f(r,X_{r}^{-k\tau}){\,\rm{d}}{r}\Big{|}^{q}\bigg{]}+\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}g(r,X_{r}^{-k\tau}){\,\rm{d}}{W_{r}}\Big{|}^{q}\bigg{]}.$
By the Hölder inequlity, (2.8) and Lemma 2.2, we have
$~{}\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}-AX_{r}^{-k\tau}+f(r,X_{r}^{-k\tau}){\,\rm{d}}{r}\Big{|}^{q}\bigg{]}$
$\leq$ $~{}(t-s)^{q-1}\int_{s}^{t}\mathbb{E}\big{[}|-AX_{r}^{-k\tau}+f(r,X_{r}^{-k\tau})|^{q}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{q-1}\int_{s}^{t}\mathbb{E}\big{[}|X_{r}^{-k\tau}|^{q}\big{]}+\mathbb{E}\big{[}|f(r,X_{r}^{-k\tau})|^{q}\big{]}{\,\rm{d}}{r}$ (4.3)
$\leq$ $~{}C(t-s)^{q-1}\int_{s}^{t}1+\mathbb{E}\big{[}|X_{r}^{-k\tau}|^{q}\big{]}+\mathbb{E}\big{[}|X_{r}^{-k\tau}|^{q\gamma}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{q}.$
Applying [14, Theorem 7.1 in Chapter 1], (2.8) and Lemma 2.2 yields
$\mathbb{E}\bigg{[}\Big{|}\int_{s}^{t}g(r,X_{r}^{-k\tau}){\,\rm{d}}{W_{r}}\Big{|}^{q}\bigg{]}\leq$ $~{}C(t-s)^{\frac{q}{2}-1}\int_{s}^{t}\mathbb{E}\big{[}|g(r,X_{r}^{-k\tau})|^{q}\big{]}{\,\rm{d}}{r}$
$\leq$ $~{}C(t-s)^{\frac{q}{2}-1}\int_{s}^{t}1+\mathbb{E}\big{[}|X_{r}^{-k\tau}|^{\frac{q(\gamma+1)}{2}}\big{]}{\,\rm{d}}{r}$ (4.4)
$\leq$ $~{}C(t-s)^{\frac{q}{2}}.$
Substituting (4) and (4) into (4) enables us to derive (4.1). ∎
Next, we define
$\mathcal{R}_{k,j}:=$ $~{}\theta\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j+1}}^{-k\tau}\big{)}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})\big{)}{\,\rm{d}}{s}$
$~{}+(1-\theta)\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j}}^{-k\tau}\big{)}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j},X_{-k\tau+t_{j}}^{-k\tau})\big{)}{\,\rm{d}}{s}$ (4.5)
$~{}+\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}g(s,X_{s}^{-k\tau})-g(t_{j},X_{-k\tau+t_{j}}^{-k\tau}){\,\rm{d}}{W_{s}}$
for all $k,j\in\mathbb{N}$. The following result provides uniform bounded estimates for the second moment of $\mathcal{R}_{k,j}$ and its conditional expectation $\mathbb{E}(\mathcal{R}_{k,j}~{}|~{}\mathcal{F}_{-k\tau+t_{j}})$.
Lemma 4.2.
Suppose that Assumption 2.1 holds. Then for any $k,j\in\mathbb{N}$, there exists a constant $C>0$, independent of $k,j$, such that
$\mathbb{E}\big{[}|\mathcal{R}_{k,j}|^{2}\big{]}\leq C\Delta^{2},\quad\mathbb{E}\big{[}|\mathbb{E}(\mathcal{R}_{k,j}~{}|~{}\mathcal{F}_{-k\tau+t_{j}})|^{2}\big{]}\leq C\Delta^{3}.$ (4.6)
Proof.
By the Hölder inequality and the Itô isometry, we obtain
$\mathbb{E}\big{[}|\mathcal{R}_{k,j}|^{2}\big{]}\leq$ $~{}3\theta^{2}\mathbb{E}\bigg{[}\Big{|}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j+1}}^{-k\tau}\big{)}$
$~{}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})\big{)}{\,\rm{d}}{s}\Big{|}^{2}\bigg{]}$
$~{}+3(1-\theta)^{2}\mathbb{E}\bigg{[}\Big{|}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j}}^{-k\tau}\big{)}$
$~{}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j},X_{-k\tau+t_{j}}^{-k\tau})\big{)}{\,\rm{d}}{s}\Big{|}^{2}\bigg{]}$
$~{}+3\mathbb{E}\bigg{[}\Big{|}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}g(s,X_{s}^{-k\tau})-g(t_{j},X_{-k\tau+t_{j}}^{-k\tau}){\,\rm{d}}{W_{s}}\Big{|}^{2}\bigg{]}$ (4.7)
$\leq$ $~{}C\Delta\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}\mathbb{E}\big{[}|(-A)(X_{s}^{-k\tau}-X_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$~{}+\mathbb{E}\big{[}|(f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}{\,\rm{d}}{s}$
$~{}+C\Delta\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}\mathbb{E}\big{[}|(-A)(X_{s}^{-k\tau}-X_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}$
$~{}+\mathbb{E}\big{[}|f(s,X_{s}^{-k\tau})-f(t_{j},X_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}{\,\rm{d}}{s}$
$~{}+C\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}\mathbb{E}\big{[}|g(s,X_{s}^{-k\tau})-g(t_{j},X_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}{\,\rm{d}}{s}.$
Now for any $s\in[-k\tau+t_{j},-k\tau+t_{j+1}]$, using (2.3), (2.7) and Lemma 4.1 indicates
$~{}\mathbb{E}\big{[}|(f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|(f(s,X_{s}^{-k\tau})-f(-k\tau+t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})|^{2}\big{]}$
$\leq$ $~{}C\mathbb{E}\big{[}\big{(}1+|X_{s}^{-k\tau}|+|X_{-k\tau+t_{j+1}}^{-k\tau}|\big{)}^{2(\gamma-1)}|X_{s}^{-k\tau}-X_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+C\mathbb{E}\big{[}\big{(}1+|X_{s}^{-k\tau}|+|X_{-k\tau+t_{j+1}}^{-k\tau}|\big{)}^{2\gamma}|s-(-k\tau+t_{j+1})|^{2}\big{]}$ (4.8)
$\leq$ $~{}C\big{(}\mathbb{E}\big{[}\big{(}1+|X_{s}^{-k\tau}|+|X_{-k\tau+t_{j+1}}^{-k\tau}|\big{)}^{4\gamma-2}\big{]}\big{)}^{\frac{2(\gamma-1)}{4\gamma-2}}$
×$~{}\times\big{(}\mathbb{E}\big{[}|X_{s}^{-k\tau}-X_{-k\tau+t_{j+1}}^{-k\tau}|^{\frac{4\gamma-2}{\gamma}}\big{]}\big{)}^{\frac{2\gamma}{4\gamma-2}}$
$~{}+C\Delta^{2}\mathbb{E}\big{[}\big{(}1+|X_{s}^{-k\tau}|+|X_{-k\tau+t_{j+1}}^{-k\tau}|\big{)}^{2\gamma}\big{]}$
$\leq$ $C\Delta,$
and similarly
$\mathbb{E}\big{[}|f(s,X_{s}^{-k\tau})-f(t_{j},X_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}\leq C\Delta,$ (4.9)
$\mathbb{E}\big{[}|g(s,X_{s}^{-k\tau})-g(t_{j},X_{-k\tau+t_{j}}^{-k\tau})|^{2}\big{]}\leq C\Delta.$ (4.10)
Plugging (4), (4.9) and (4.10) into (4) promises the first part of (4.6). For the second part of (4.6), we use the Jensen inequality for conditional expectation to get
$~{}\mathbb{E}\big{[}|\mathbb{E}(\mathcal{R}_{k,j}~{}|~{}\mathcal{F}_{-k\tau+t_{j}})|^{2}\big{]}$
$\leq$ $~{}2\theta^{2}\mathbb{E}\bigg{[}\Big{|}\mathbb{E}\Big{(}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j+1}}^{-k\tau}\big{)}$
$~{}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})\big{)}{\,\rm{d}}{s}~{}|~{}\mathcal{F}_{-k\tau+t_{j}}\Big{)}\Big{|}^{2}\bigg{]}$
$~{}+2(1-\theta)^{2}\mathbb{E}\bigg{[}\Big{|}\mathbb{E}\Big{(}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j}}^{-k\tau}\big{)}$
$~{}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j},X_{-k\tau+t_{j}}^{-k\tau})\big{)}{\,\rm{d}}{s}~{}|~{}\mathcal{F}_{-k\tau+t_{j}}\Big{)}\Big{|}^{2}\bigg{]}$
$\leq$ $~{}2\mathbb{E}\bigg{[}\mathbb{E}\Big{(}\Big{|}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j+1}}^{-k\tau}\big{)}$
$~{}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})\big{)}{\,\rm{d}}{s}\Big{|}^{2}~{}|~{}\mathcal{F}_{-k\tau+t_{j}}\Big{)}\bigg{]}$
$~{}+2\mathbb{E}\bigg{[}\mathbb{E}\Big{(}\Big{|}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j}}^{-k\tau}\big{)}$
$~{}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j},X_{-k\tau+t_{j}}^{-k\tau})\big{)}{\,\rm{d}}{s}\Big{|}^{2}~{}|~{}\mathcal{F}_{-k\tau+t_{j}}\Big{)}\bigg{]}$
$=$ $~{}2\mathbb{E}\bigg{[}\Big{|}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j+1}}^{-k\tau}\big{)}$
$~{}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j+1},X_{-k\tau+t_{j+1}}^{-k\tau})\big{)}{\,\rm{d}}{s}\Big{|}^{2}\bigg{]}$
$~{}+2\mathbb{E}\bigg{[}\Big{|}\int_{-k\tau+t_{j}}^{-k\tau+t_{j+1}}(-A)\big{(}X_{s}^{-k\tau}-X_{-k\tau+t_{j}}^{-k\tau}\big{)}$
$~{}+\big{(}f(s,X_{s}^{-k\tau})-f(t_{j},X_{-k\tau+t_{j}}^{-k\tau})\big{)}{\,\rm{d}}{s}\Big{|}^{2}\bigg{]}.$
Repeating the techniques used in (4) and exploiting (4), (4.9) as well as (4.10) implies the second part of (4.6). Thus we complete the proof. ∎
It is expected that the orders of estimates with respect to the stepsize $\Delta$ in (4.6) should be higher when SDEs (1.1) driven by additive noise. We will show that this happens at the price of requiring a differentiable condition on the drift coefficient. Following the idea presented in [7], we make the following assumption.
Assumption 4.3.
Suppose that the diffusion coefficient function $g\colon\mathbb{R}\to\mathbb{R}^{d\times m}$ is continuous and periodic in time with period $\tau>0$, i.e., $g(t+\tau)=g(t)$ for all $t\in\mathbb{R}$. Besides, there exists a constant $C>0$ such that $\sup\limits_{t\in[0,\tau)}|g(t)|\leq C$ and
$|g(t)-g(s)|\leq C|t-s|,\quad s,t\in[0,\tau).$ (4.11)
Moreover, assume that the drift coefficient function $f\colon\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ is continuously differentiable, and that there exists a constant $C>0$ such that
$\bigg{|}\bigg{(}\frac{\partial{f}(t,y)}{\partial{x}}-\frac{\partial{f}(t,z)}{\partial{x}}\bigg{)}u\bigg{|}\leq C(1+|y|+|z|)^{\gamma-2}|y-z||u|,\quad y,z,u\in\mathbb{R}^{d},$ (4.12)
where $\gamma\geq 1$ comes from (2.3).
Under the above additional assumption, one can repeat the arguments used in [7, Theorem 4.6] to improve the estimates in Lemma 4.2. The proof of the following lemma is thus omitted.
Lemma 4.4.
Suppose that Assumptions 2.1 and 4.3 hold. Then for any $k,j\in\mathbb{N}$, there exists a constant $C>0$, independent of $k,j$, such that
$\mathbb{E}\big{[}|\mathcal{R}_{k,j}|^{2}\big{]}\leq C\Delta^{3},\quad\mathbb{E}\big{[}|\mathbb{E}(\mathcal{R}_{k,j}~{}|~{}\mathcal{F}_{-k\tau+t_{j}})|^{2}\big{]}\leq C\Delta^{4}.$ (4.13)
Now we are in a position to present a uniform estimate for the error between the numerical solution $\hat{X}_{-k\tau+t_{j}}^{-k\tau}$ and the exact solution $X_{-k\tau+t_{j}}^{-k\tau}$ in the mean square sense.
Lemma 4.5.
Suppose that Assumption 2.1 holds. Then there exists a constant $C>0$, independent of $j,k$ and $\Delta$, such that
$\sup_{k,j\in\mathbb{N}}|X_{-k\tau+t_{j}}^{-k\tau}-\hat{X}_{-k\tau+t_{j}}^{-k\tau}|_{L^{2}(\Omega;\mathbb{R}^{d})}\leq C\Delta^{\frac{1}{2}}.$ (4.14)
If in addition Assumption 4.3 holds, then there exists a constant $C>0$, independent of $j,k$ and $\Delta$, such that
$\sup_{k,j\in\mathbb{N}}|X_{-k\tau+t_{j}}^{-k\tau}-\hat{X}_{-k\tau+t_{j}}^{-k\tau}|_{L^{2}(\Omega;\mathbb{R}^{d})}\leq C\Delta.$ (4.15)
Proof.
For any $k,j\in\mathbb{N}$ and $t\in\mathbb{R}$, let us first denote
$f_{-k\tau+t}^{-k\tau}:=f(t,X_{-k\tau+t}^{-k\tau}),\qquad\hat{f}_{-k\tau+t_{j}}^{-k\tau}:=f(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau}),$
$g_{-k\tau+t}^{-k\tau}:=g(t,X_{-k\tau+t}^{-k\tau}),\qquad\hat{g}_{-k\tau+t_{j}}^{-k\tau}:=g(t_{j},\hat{X}_{-k\tau+t_{j}}^{-k\tau})$
and $e_{-k\tau+t_{j}}^{-k\tau}:=X_{-k\tau+t_{j}}^{-k\tau}-\hat{X}_{-k\tau+t_{j}}^{-k\tau}$. It follows from (1.1), (1) and (4) that
$~{}e_{-k\tau+t_{j+1}}^{-k\tau}-\theta\Delta\big{(}(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}\big{)}$
$=$ $~{}e_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}\big{)}$
$~{}+\big{(}g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}\big{)}\Delta W_{-k\tau+t_{j}}^{-k\tau}+\mathcal{R}_{k,j}.$
Due to that the terms $e_{-k\tau+t_{j}}^{-k\tau}$, $f_{-k\tau+t_{j}}^{-k\tau}$, $\hat{f}_{-k\tau+t_{j}}^{-k\tau}$, $g_{-k\tau+t_{j}}^{-k\tau}$ and $\hat{g}_{-k\tau+t_{j}}^{-k\tau}$ are $\mathcal{F}_{-k\tau+t_{j}}$-measurable, we apply the independence between $g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}$ and $\mathcal{F}_{-k\tau+t_{j}}$ to show
$~{}\mathbb{E}\big{[}\big{|}e_{-k\tau+t_{j+1}}^{-k\tau}-\theta\Delta\big{(}(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}\big{)}\big{|}^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}\big{|}e_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}\big{)}\big{|}^{2}\big{]}+\mathbb{E}\big{[}|\mathcal{R}_{k,j}|^{2}\big{]}$
$~{}+\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+2\mathbb{E}\big{[}\big{\langle}\big{(}g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}\big{)}\Delta W_{-k\tau+t_{j}}^{-k\tau},\mathcal{R}_{k,j}\big{\rangle}\big{]}$ (4.16)
$~{}+2\mathbb{E}\big{[}\big{\langle}e_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}\big{)},\mathcal{R}_{k,j}\big{\rangle}\big{]}.$
As a consequence of (2.2) and (2.4), we obtain
$~{}\mathbb{E}\big{[}\big{|}e_{-k\tau+t_{j+1}}^{-k\tau}-\theta\Delta\big{(}(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}\big{)}\big{|}^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|e_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+2\theta\Delta\mathbb{E}\big{[}\langle e_{-k\tau+t_{j+1}}^{-k\tau},Ae_{-k\tau+t_{j+1}}^{-k\tau}\rangle\big{]}-2\theta\Delta\mathbb{E}\big{[}\langle e_{-k\tau+t_{j+1}}^{-k\tau},f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}\rangle\big{]}$ (4.17)
$\geq$ $~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j+1}}^{-k\tau}-\hat{g}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]},$
and
$~{}\mathbb{E}\big{[}\big{|}e_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}\big{)}\big{|}^{2}\big{]}$
$=$ $~{}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+2(1-\theta)\Delta\mathbb{E}\big{[}\big{\langle}e_{-k\tau+t_{j}}^{-k\tau},(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}\big{\rangle}\big{]}$ (4.18)
$\leq$ $~{}\big{(}1-2(\lambda-L_{f})(1-\theta)\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}-2(p^{*}-1)(1-\theta)\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}.$
Utilizing the weighted Young inequality $2ab\leq\kappa a^{2}+\frac{b^{2}}{\kappa},a,b\in\mathbb{R}$ with $\kappa:=\min\{\lambda-L_{f},2\theta-1\}>0$ leads to
$~{}2\mathbb{E}\big{[}\big{\langle}\big{(}g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}\big{)}\Delta W_{-k\tau+t_{j}}^{-k\tau},\mathcal{R}_{j}\big{\rangle}\big{]}$
$\leq$ $~{}(p^{*}-2)\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+\frac{1}{p^{*}-2}\mathbb{E}\big{[}|\mathcal{R}_{k,j}|^{2}\big{]},$ (4.19)
and
$~{}2\mathbb{E}\big{[}\big{\langle}e_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}\big{)},\mathcal{R}_{k,j}\big{\rangle}\big{]}$
$=$ $~{}2\mathbb{E}\big{[}\big{\langle}e_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}\big{)},\mathbb{E}(\mathcal{R}_{k,j}~{}|~{}\mathcal{F}_{-k\tau+t_{j}})\big{\rangle}\big{]}$ (4.20)
$\leq$ $~{}\kappa\Delta\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}+(1-\theta)\Delta\big{(}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}\big{)}|^{2}\big{]}$
$~{}+\frac{1}{\kappa\Delta}\mathbb{E}\big{[}|\mathbb{E}(\mathcal{R}_{k,j}~{}|~{}\mathcal{F}_{-k\tau+t_{j}})|^{2}\big{]}.$
Substituting (4), (4), (4) and (4) into (4) yields
$~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j+1}}^{-k\tau}-\hat{g}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$\leq$ $~{}(1+\kappa\Delta)\big{(}1-2(\lambda-L_{f})(1-\theta)\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$ (4.21)
$~{}+\big{(}1-2(1+\kappa\Delta)(1-\theta)\big{)}(p^{*}-1)\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+(1+\kappa\Delta)(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+\frac{p^{*}-1}{p^{*}-2}\mathbb{E}\big{[}|\mathcal{R}_{k,j}|^{2}\big{]}+\frac{1}{\kappa\Delta}\mathbb{E}\big{[}|\mathbb{E}(\mathcal{R}_{k,j}~{}|~{}\mathcal{F}_{-k\tau+t_{j}})|^{2}\big{]}.$
By means of Lemma 4.2, we obtain
$~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j+1}}^{-k\tau}-\hat{g}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$\leq$ $~{}(1+\kappa\Delta)\big{(}1-2(\lambda-L_{f})(1-\theta)\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$ (4.22)
$~{}+\big{(}1-2(1+\kappa\Delta)(1-\theta)\big{)}(p^{*}-1)\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+(1+\kappa\Delta)(1-\theta)^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+C\Delta^{2}.$
From $\kappa=\min\{\lambda-L_{f},2\theta-1\}>0$ and $2(1-\theta)^{2}\Delta\leq 1$, one gets
$~{}(1+\kappa\Delta)\big{(}1-2(\lambda-L_{f})(1-\theta)\Delta\big{)}$
$=$ $~{}1+\kappa\Delta-2(\lambda-L_{f})\Delta+2(\lambda-L_{f})\theta\Delta-2\kappa(\lambda-L_{f})(1-\theta)\Delta^{2}$
$\leq$ $~{}1+2(\lambda-L_{f})\theta\Delta-(\lambda-L_{f})\Delta$
$=$ $~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\bigg{(}1-\frac{(\lambda-L_{f})\Delta}{1+2(\lambda-L_{f})\theta\Delta}\bigg{)},$
as well as
$1-2(1+\kappa\Delta)(1-\theta)=-1+2\theta-2\kappa\Delta(1-\theta)\leq 2\theta-1=2\theta\bigg{(}1-\frac{1}{2\theta}\bigg{)},$
and
$(1+\kappa\Delta)(1-\theta)^{2}=$ $~{}\theta^{2}-(2\theta-1)+\kappa\Delta(1-\theta)^{2}$
$\leq$ $~{}\theta^{2}-(2\theta-1)\big{(}1-\Delta(1-\theta)^{2}\big{)}$
$\leq$ $~{}\theta^{2}-\frac{2\theta-1}{2}=\theta^{2}\bigg{(}1-\frac{2\theta-1}{2\theta^{2}}\bigg{)}.$
Setting
$\overline{C_{\Delta}}:=\max\bigg{\{}1-\frac{(\lambda-L_{f})\Delta}{1+2(\lambda-L_{f})\theta\Delta},~{}1-\frac{1}{2\theta},~{}1-\frac{2\theta-1}{2\theta^{2}}\bigg{\}}\in(0,1)$ (4.23)
and employing (4) indicate
$~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j+1}}^{-k\tau}-\hat{g}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$\leq$ $~{}\overline{C_{\Delta}}\Big{(}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}\Big{)}+C\Delta^{2},$
which implies
$~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j+1}}^{-k\tau}-\hat{g}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j+1}}^{-k\tau}+f_{-k\tau+t_{j+1}}^{-k\tau}-\hat{f}_{-k\tau+t_{j+1}}^{-k\tau}|^{2}\big{]}-\frac{C\Delta^{2}}{1-\overline{C_{\Delta}}}$
$\leq$ $~{}\overline{C_{\Delta}}\bigg{(}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}-\frac{C\Delta^{2}}{1-\overline{C_{\Delta}}}\bigg{)}$
It follows from an iterative argument and $e_{-k\tau}^{-k\tau}=0$, $f_{-k\tau}^{-k\tau}=\hat{f}_{-k\tau}^{-k\tau}$, $g_{-k\tau}^{-k\tau}=\hat{g}_{-k\tau}^{-k\tau}$ that
$~{}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau+t_{j}}^{-k\tau}-\hat{g}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau+t_{j}}^{-k\tau}+f_{-k\tau+t_{j}}^{-k\tau}-\hat{f}_{-k\tau+t_{j}}^{-k\tau}|^{2}\big{]}-\frac{C\Delta^{2}}{1-\overline{C_{\Delta}}}$
$\leq$ $~{}\overline{C_{\Delta}}^{j}\bigg{(}\big{(}1+2(\lambda-L_{f})\theta\Delta\big{)}\mathbb{E}\big{[}|e_{-k\tau}^{-k\tau}|^{2}\big{]}+2(p^{*}-1)\theta\Delta\mathbb{E}\big{[}|g_{-k\tau}^{-k\tau}-\hat{g}_{-k\tau}^{-k\tau}|^{2}\big{]}$
$~{}+\theta^{2}\Delta^{2}\mathbb{E}\big{[}|(-A)e_{-k\tau}^{-k\tau}+f_{-k\tau}^{-k\tau}-\hat{f}_{-k\tau}^{-k\tau}|^{2}\big{]}-\frac{C\Delta^{2}}{1-\overline{C_{\Delta}}}\bigg{)}$
$\leq$ $~{}0$
Together with
$\frac{1}{1-\overline{C_{\Delta}}}=\max\bigg{\{}\frac{1+2(\lambda-L_{f})\theta\Delta}{(\lambda-L_{f})\Delta},~{}2\theta,~{}\frac{2\theta^{2}}{2\theta-1}\bigg{\}}\leq\Delta^{-1}\max\bigg{\{}\frac{1+2(\lambda-L_{f})\theta}{\lambda-L_{f}},~{}2\theta,~{}\frac{2\theta^{2}}{2\theta-1}\bigg{\}},$
we derive the desired result (4.14). Concerning (4.15), it suffices to put (4.13) into (4) and continue the remainder steps as that to derive (4.14). Thus we complete the proof. ∎
Armed with the above preparations, we are able to reveal the main result of this work.
Theorem 4.6.
Let $X_{t}^{*}$ be the random periodic solution of SDE (1.1) and $\hat{X}_{t}^{*}$ the random periodic solution of the stochastic theta approximation (1). Suppose that Assumption 2.1 holds. Then there exists $C>0$, independent of $\Delta$, such that
$|X_{t}^{*}-\hat{X}_{t}^{*}|_{L^{2}(\Omega;\mathbb{R}^{d})}\leq C\Delta^{\frac{1}{2}}.$ (4.24)
If in addition Assumption 4.3 holds, then there exists $C>0$, independent of $\Delta$, such that
$|X_{t}^{*}-\hat{X}_{t}^{*}|_{L^{2}(\Omega;\mathbb{R}^{d})}\leq C\Delta.$ (4.25)
Proof.
Based on (2.12) and (3.8), we utilize Lemma 4.5 and
[…]
to obtain (4.24) and (4.25) and finish the proof. ∎
5 Numerical experiments
Some numerical experiments will be performed to illustrate the previous theoretical results in this section. Let us focus on the following one-dimensional SDE
$\text{d}X^{t_{0}}_{t}=\big{(}-\lambda X^{t_{0}}_{t}-a(X^{t_{0}}_{t})^{3}(1+\sin(\pi t))\big{)}\text{d}t+\big{(}b+cX^{t_{0}}_{t}+d(X^{t_{0}}_{t})^{2}(1+\sin(\pi t))\big{)}\text{d}W_{t}$ (5.1)
for all $t\geq t_{0}$, where $\lambda,a>0$ and $b,c,d\in\mathbb{R}$. Compared with (1.1), we know that
$A=\lambda,\quad f(t,x)=-ax^{3}(1+\sin(\pi t)),\quad g(t,x)=b+cx+dx^{2}(1+\sin(\pi t)),\quad x\in\mathbb{R},t\in\mathbb{R}.$ (5.2)
Obviously, (A1) holds. Besides, $f$ and $g$ are continuous and periodic in time with period $\tau=2>0$, i.e., (A2) holds. Noting that for any $x,y\in\mathbb{R}$ and $s,t\in\mathbb{R}$, we have
$|f(t,x)-f(s,y)|=$ $~{}|-a(x^{3}-y^{3})(1+\sin(\pi t))-ay^{3}(\sin(\pi t)-\sin(\pi s))|$
$\leq$ $~{}2a(|x|^{2}+|x||y|+|y|^{2})|x-y|+a\pi|y|^{3}|t-s|$ (5.3)
$\leq$ $~{}2a(1+|x|+|y|)^{2}|x-y|+a\pi(1+|x|+|y|)^{3}|t-s|,$
and
$~{}(x-y)(f(t,x)-f(s,y))$
$=$ $~{}(x-y)\big{(}-a(x^{3}-y^{3})(1+\sin(\pi t))-ay^{3}(\sin(\pi t)-\sin(\pi s))\big{)}$
$\leq$ $~{}-a|x-y|^{2}(x^{2}+xy+y^{2})(1+\sin(\pi t))+a\pi(|x||y|^{3}+|y|^{4})|t-s|,$
as well as
$|g(t,x)-g(s,y)|^{2}=$ $~{}|c(x-y)+d(x^{2}-y^{2})(1+\sin(\pi t))+dy^{2}(\sin(\pi t)-\sin(\pi s))|^{2}$
$\leq$ $~{}3c^{2}|x-y|^{2}+6d^{2}|x-y|^{2}(x+y)^{2}(1+\sin(\pi t))+6\pi^{2}d^{2}|y|^{4}|t-s|.$
By requiring
$12d^{2}(p^{*}-1)\leq a,$ (5.4)
one can show that for any $x,y\in\mathbb{R}$ and $s,t\in\mathbb{R}$,
$~{}(x-y)(f(t,x)-f(s,y))+(p^{*}-1)|g(t,x)-g(s,y)|^{2}$
$\leq$ $~{}\big{(}3c^{2}(p^{*}-1)-\big{(}a(x^{2}+xy+y^{2})-6d^{2}(p^{*}-1)(x+y)^{2}\big{)}(1+\sin(\pi t))\big{)}|x-y|^{2}$
$~{}+a\pi(|x||y|^{3}+|y|^{4})|t-s|+6\pi^{2}d^{2}(p^{*}-1)|y|^{4}|t-s|$
$\leq$ $~{}3c^{2}(p^{*}-1)|x-y|^{2}+\big{(}2a\pi+6\pi^{2}d^{2}(p^{*}-1)\big{)}(1+|x|^{4}+|y|^{4})|t-s|,$
which together with (5) indicates that (A3) holds with $L_{f}=3c^{2}(p^{*}-1)$ and $\gamma=3$. In combination with (5.4), $L_{f}\in(0,\lambda)$ and $p^{*}>4\gamma-2$, we choose
$p^{*}=21,\quad\lambda=5\pi,\quad a=3,\quad b=1.5,\quad c=0.5,\quad d=0.1$
such that all conditions in Assumption 2.1 are fulfilled. It follows that (5.1) admits a unique random periodic solution. Before performing experiments, we emphasize that in what follows the Newton iterations with precision $10^{-5}$ are employed to solve the nonlinear equation arising from the implementation of the implicit ST method in every time step.
(a) $\theta=0.75$
(b) $\theta=1$
Figure 1: Random periodic solution does not dependent on the initial values
(a) $\theta=0.75$
(b) $\theta=1$
Figure 2: Validate periodicity based on $\hat{X}_{t}^{*}(\Theta_{-\tau}\omega)=\hat{X}_{t-\tau}^{*}(\omega)$
(a) $\theta=0.75$
(b) $\theta=1$
Figure 3: Validate periodicity via the periodicity of $\hat{X}^{0}(t,\Theta_{-t}\omega)$
Theorem 3.3 indicates that for each $\theta\in(1/2,1]$, the corresponding ST method applied to (5.1) admits a unique random periodic solution. Let us first numerically verify that the random periodic solution of ST method does not dependent on the initial values. To this end, we choose three different initial values $\xi=\pm 0.6,0$ for $\theta=0.75,1$, the corresponding paths $\hat{X}_{t}^{-10}(\omega,0.6)$, $\hat{X}_{t}^{-10}(\omega,0)$ and $\hat{X}_{t}^{-10}(\omega,-0.6)$ are simulated on the time interval $[-10,0]$ with stepsize $\Delta=0.1$; see Figure 1. It can be seen that for each $\theta$, the three paths coincide after a very short time, which means that taking pull-back time $-10$ is sufficient to obtain a good convergence to the random periodic paths for $t\geq-8$. Following the arguments in [4], we will verify the periodicity in two ways. Based on $\hat{X}_{t}^{*}(\Theta_{-\tau}\omega)=\hat{X}_{t-\tau}^{*}(\omega)$, we will plot the paths $\hat{X}_{t}^{-10}(\omega,0.6),t\in[-4,0]$ and $\hat{X}_{t}^{-10}(\Theta_{-2}\omega,0.6),t\in[-6,-2]$ for $\theta=0.75,1$ with stepsize $\Delta=0.1$ in Figure 2, which shows that the two segmented processes resemble each other with the period $\tau=2$. At last, we check whether or not $\hat{X}_{t}^{*}(t,\Theta_{-\tau}\omega)$ is periodic with period $\tau=2$. For this purpose, the path $X^{0}(t,\Theta_{-t}\omega),t\in[0,10]$ with $X^{0}(0,\Theta_{-0}\omega)=-0.2$ is given in Figure 3 with stepsize $\Delta=0.1$. Then we obtain a periodic pull-back path with period $\tau=2$, which validates the random periodicity of the original path.
(a) $\theta=0.75$
(b) $\theta=1$
Figure 4: Mean square convergence orders for (5.1)
Theorem 4.6 shows that for every $\theta\in(1/2,1]$, the random periodic solution converges to that of (5.1) with order $1/2$ in the mean square sense. To support this result numerically, we identify the unavailable exact solution with a numerical approximation generated by the ST method with a fine stepsize $\Delta=2^{-16}$ on the time interval $[-10,10]$. The other numerical approximations are calculated with five different equidistant stepsize $\Delta=2^{-i},i=10,\cdots,14$. Here the expectations are approximated by the Monte Carlo simulation with 500 different Brownian paths. Figure 4 shows that for different $\theta$, the root mean square error line and the reference line appear to parallel to each other, indicating that the mean square convergence rate of the ST method is $1/2$. A least square fit indicates that the slope of the line for the ST method is $0.5804$ for $\theta=0.75$ and $0.6208$ for $\theta=1$, identifying the theoretical result in Theorem 4.6.
(a) $\theta=0.75$
(b) $\theta=1$
Figure 5: Mean square convergence orders for (5.5)
For the additive case, we consider the following one-dimensional SDE
${\,\rm{d}}{X_{t}^{t_{0}}}=-10\pi X_{t}^{t_{0}}{\,\rm{d}}{t}+\sin(2\pi t){\,\rm{d}}{t}+0.05{\,\rm{d}}{W_{t}},\quad t\geq t_{0}.$ (5.5)
One can check that the associated period is $1$ and Assumptions 2.1 and 4.3 are fulfilled. It is worth noting that (5.5) has been considered in [22] to test the periodicity and mean square convergence of the backward Euler method. Thus for simplicity, we only verify the mean square order of ST methods on the time interval $[-10,10]$ with $\theta=0.75,1$ under the above settings for convergence test of (5.1). Figure 5 shows that the slopes of the error lines and the reference lines match well, indicating that the ST methods have strong rates of order one order one in additive case. Actually, least square fits produce rate $1.0289$ for $\theta=0.75$ and rate $1.0289$ for $\theta=1$. These facts coincide with the previous theoretical result in Theorem 4.6.